(40) Use Cramers rule to solve for z: x+2y-3z=-24
2x-y+3z=13
x+2y+3z=0
POssible Answers:
(A) -5
(B) -6
(C) -2
(D) 4
Thanks for your help

Answer by mukhopadhyay(490)

Determinat D = Determinant of [1,2,-3;2,-1,3;1,2,3]
= 1[(-1)(3)-(2*3)] - 2[2*3-2*(-3)] + 1(2*3-(-1)(-3)]
= 1*(-9) - 2*12 + 3
=-9-24+3 = -30
D(z) = Determinant of [1,2,-24;2,-1,13;1,2,0]
= 1(-26) - 2(48) + 1(26-24)
= -26 - 96 + 2
= -122 + 2
= -120
z = D(z)/D = -120/-30 = 4
Correction option is D